OpenStudy (anonymous):

Solve for x>0, e^x^e^x^e^x^... = 2

OpenStudy (anonymous):

\[e^{x^{e^{x}}} \] till infinity =2 = y (say) \[=e^{x^{y}}\] \[=e^{x^{2}}\] =2 Therefore, \[x^2 = \ln2\] x = ln2 Greetings!

OpenStudy (anonymous):

Oh wow, thanks, I didn't think about that approach!

OpenStudy (anonymous):

welcome :-)

OpenStudy (anonymous):

Sorry, I know its obvious but its \[\sqrt{\ln2}\]